# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ A.3 Guided Diffusion Models
Guided diffusion aims to sample from the noisy data distribution conditioned on some property $y$, $\mathbf{x}\sim p(\mathbf{x}_{0}|\mathbf{z}_{T},y)$, such that $p(y|\mathbf{x}_{0})$ is maximized. Therefore, the marginal reverse transition aims to sample from a distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0},y)$ conditioned on the target unmasked sequence $\mathbf{x}_{0}$ and a property $y$. Since we aim to reconstruct the sequence $\mathbf{x}_{0}$ in the reverse process, the learned distribution is conditioned on a learned unmasked sequence $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$. Using Baye’s theorem, we can decompose the guided conditional distribution as:
@@
Classifier-Free Guidance. Classifier-free guidance strategies aim to model the conditional distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},y)$ by directly training the diffusion model on directly on a subset of the unconditional data with property $y$, such that after training, the model samples from a learned distribution $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t},y)$. However, classifier-free guidance fails at tasks where quality annotated data is scarce, including peptide generation. Furthermore, this strategy cannot scale to multiple objectives which would require data conditioned on more than one property.
@@
$p_{\theta,\phi}(\mathbf{z}_{s}|\mathbf{z}_{t},y)=\frac{p_{\phi}(y|\mathbf{z}_{s})p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})}{p(y|\mathbf{z}_{t})}$ (50)
Since the model parameters implicitly learn the normalized distribution, we can drop the $p(y|\mathbf{z}_{t})$ term. Then, at each iteration, we update the parameters $\theta,\phi$ in the direction of the gradient of $\log p_{\theta,\phi}(\mathbf{z}_{s}|\mathbf{z}_{t},y)$ obtained by combining the unconditional distribution $\log p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$ and classifier probability $p_{\phi}(y|\mathbf{z}_{s})$ with respect to the sampled sequence $\mathbf{z}_{s}$.
@@
Unlike classifier-free guidance, classifier-based guidance does not require training a generative model from a conditioned dataset. However, the gradient-based strategy for classifier-based guidance is not directly applicable to discrete state spaces due to the lack of a defined gradient. To mimic gradient-based updates to each sampling step, Gruver et al. [111] leveraged iterative guidance steps on hidden latent embeddings $\mathbf{h}_{t}$ for each token before projecting back to a discrete categorical distribution $p_{\theta}(\mathbf{z}_{s}|\mathbf{h}_{t})$ at each time step. However, projecting to and from the continuous and discrete spaces results in inconsistencies in the guided diffusion process, where optimized hidden embeddings do not always map to optimal tokens.
@@
However, computing $p_{\phi}(y|\mathbf{z}^{\prime}_{s})$ for all the possible transitions from state $\mathbf{z}_{t}$ is infeasible. Previous work has bypassed this limitation by approximation. For continuous and differentiable classifier functions $p(y|\mathbf{x}):\mathbb{R}^{L\times|\mathcal{V}|}\to\mathbb{R}$, we can approximate the denominator using the first-order Taylor expansion given by
$\log p_{\phi}(y|\mathbf{z}_{s},\mathbf{z}_{t})$ $\approx\log p_{\phi}(y|\mathbf{z}_{t})+(\mathbf{x}_{s}-\mathbf{x}_{t})^{\top}\nabla_{\mathbf{z}}\log p(y|\mathbf{z})|_{\mathbf{z}=\mathbf{z}_{t}}$ (53)
which approximates the likelihood of observing property $p$ at the slightly denoised state $\mathbf{z}_{s}=\mathbf{z}_{t-\frac{1}{T}}$ given the known log-probability of observing $y$ for state $\mathbf{x}_{t}$. This eliminates the need to explicitly sample $\mathbf{x}_{s}$ for all possible state transitions and compute $\log p_{\phi}(y|\mathbf{z}_{s},\mathbf{z}_{t})$.
This approximation strategy has been used by Digress [112] for guided discrete diffusion on categorical graph generation. Furthermore, Nisonoff et al. [27] uses the Taylor-approximated conditional distribution $\log p_{\phi}(y|\mathbf{z}_{s})$ to adjust the unconditional transition rates $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s}|y)$ given the unconditional rates $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s})$ for predictor-guidance of Continuous-Time Markov Chains (CTMCs) in the discrete state space.
@@
where $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s}|y)$ is the predictor-adjusted rate of transitioning from state $\mathbf{z}_{t}$ to state $\mathbf{z}_{t}$
This strategy fails to scale to multiple objective guidance since it would require computing the joint probability over $K$ objectives $p_{\phi}(y_{1},y_{2},\dots,y_{K}|\mathbf{z}_{s},\mathbf{z}_{t})$ for some $K>1$. If all properties are mutually independent, we can factorize the distribution and compute the estimated probability of each objective and take their product $p_{\phi}(y_{1},y_{2},\dots,y_{K}|\mathbf{z}_{s},\mathbf{z}_{t})=\prod_{k=1}^{K}p_{\phi}(y_{k}|\mathbf{z}_{s},\mathbf{z}_{t})$. For the majority of multi-objective tasks including therapeutic peptide generation, independence across properties is not a reasonable assumption, and computing the joint distribution is required. Moreover, for objectives that guide toward contradictory optimal rates or transitions, training a model conditioned on these objectives could prevent the model from generating optimal sequences for either objective. Given these limitations, there remains a gap for efficient classifier-based conditioning for discrete diffusion that is robust to multi-objective tasks, which we address in this work.
@@ B.1 State-Dependent Masking Schedule
From Equation (37), we define the continuous-time forward masking probability $1-\alpha_{t}$ at time $t$ with $\alpha_{t}=\exp(-\sigma(t))$ , where $\sigma(t):[0,1]\to\mathbb{R}^{+}$. Following Lou et al. [113], we apply a log-linear masking schedule $\sigma(t)=-\log(1-t)$ for the forward diffusion process which has shown to result in the lowest variance in the negative likelihood bound (NELBO) loss [103]. Therefore, the continuous-time probability of remaining unmasked at time $t$ is equal to $\alpha_{t}=\exp{-(-\log(1-t))}=1-t$ and the weight that scales the negative log loss (NLL) is given by $\frac{1}{t}$ by our derivation in Appendix B.3.
For peptide-bond tokens, we alter the masking schedule such that peptide-bonds are masked at slower rate at earlier time steps by defining a log-polynomial masking schedule $\sigma(t)=-\log(1-t^{w})$, for some constant exponent $w\geq 1$. Note that when $w=1$, the log-polynomial schedule reduces to the log-linear schedule. Therefore, the probability of remaining unmasked becomes $\alpha_{t}=\exp\big{(}\log(1-t^{w})\big{)}=1-t^{w}$ and the weight that scales the negative log loss (NLL) is given by $\frac{w}{t}$ by our derivation in Appendix B.3.
@@
This means that the NLL of the peptide-bond tokens is weighted more heavily than non-peptide bond tokens for $t$ close to 1. As $t\to 0$, the NLL weight approaches $\infty$ for all tokens. This biases the reverse diffusion process to unmask peptide bond tokens earlier since it was trained to minimize the loss associated with each unmasking step. As $t\to 0$, the large NLL weight ensures that the final unmasking steps during the reverse diffusion process result in an unmasked sequence that lies within the space of valid peptide SMILES.
Figure 15: Plots of state-dependent masking schedules. (A) The probability of remaining unmasked during the continuous-time forward diffusion process over time $t$ given different values of $w$ as the exponent of the masking schedule $\alpha_{t}=1-t^{w}$. We use $w=1$ for non-peptide bond tokens and $w=3$ for peptide bond tokens, enabling slower masking of peptide bond tokens. (B) The weight of the negative log-loss for different exponents $w$ in the log-polynomial masking schedule. The weight of the loss is higher for larger $w$ in earlier time steps, which results in a higher penalty for wrong predictions of peptide bond tokens compared to other tokens.
@@ B.2 Derivation of State-Dependent Reverse Posterior
For a single token, the state-dependent forward diffusion process is defined by the probability distribution $q(\mathbf{z}_{t}|\mathbf{x}_{0})$ which transforms the clean inputs to sequences with varying degrees of masking based on a probability distribution given by the function $\alpha_{t}(\mathbf{x}_{0}):\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}$, which takes the token encoding $\mathbf{x}_{0}$ and outputs the probability of remaining unmasked at time $t$ depending on whether $\mathbf{x}_{0}$ encodes a peptide bond token.
@@
In this work, we classify each token into one of two states: a peptide-bond token, and all remaining tokens representing peptide side-chains and modifications. We define a function that generates a mask with values of 1 indicating tokens containing or contained within a peptide bond, and 0 otherwise (Appendix 6). Letting $\mathbf{b}\in\mathbb{R}^{|\mathcal{V}|}$ denote a vector with ones at indices of peptide bond tokens, we have $\mathbf{b}^{\top}\mathbf{x}_{0}^{(\ell)}$ when a token at position $\ell$ is a peptide bond token. Note that $\mathbf{b}$ is defined differently depending on the context of token in the full sequence which is handled by the BondMask function, and we use $\mathbf{b}$ for derivation purposes only. Then, we have $\vec{\alpha}_{t}=\big{(}\mathbf{1}-\mathbf{b}^{\top}\mathbf{x}_{0}\big{)}(1-t)+\mathbf{b}^{\top}\mathbf{x}_{0}(1-t^{w})$ or equivalently we can write
@@ B.3 Derivation of State-Dependent NELBO Loss
$\sum_{t=1}^{T-1}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}\text{KL}\bigg{(}q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})\big{|}\big{|}p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})\bigg{)}\bigg{]}$ (71)
@@
$\mathcal{L}_{\text{NELBO}}=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}T\cdot\text{KL}\bigg{(}q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})\big{|}\big{|}p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})\bigg{)}\bigg{]}$
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}T\cdot\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{w}{tT}-o\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$
@@
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{wT}{tT}-To\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$
@@
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{w}{t}-To\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$

### v2

@@ A.3 Guided Diffusion Models
Guided diffusion aims to sample from the data distribution conditioned on some property $y$, $\mathbf{x}\sim q(\mathbf{x}_{0},y)$, such that $q(y|\mathbf{x})$ is maximized. Therefore, the marginal reverse transition aims to sample $\mathbf{z}_{s}$ from a distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},y)$ conditioned on the current sequence $\mathbf{z}_{t}$ and a property $y$. Using Baye’s theorem, we can decompose the guided conditional distribution as
@@
Classifier-Free Guidance. Classifier-free guidance strategies aim to model the conditional distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},y)$ by directly training the diffusion model on a subset of the unconditional data with property $y$, such that after training, the model samples from a learned distribution $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t},y)$. However, classifier-free guidance fails at tasks where quality annotated data is scarce, including peptide sequences. Furthermore, this strategy cannot scale to multiple objectives which would require data conditioned on more than one property.
@@
$p_{\theta,\phi}(\mathbf{z}_{s}|\mathbf{z}_{t},y)=\frac{p_{\phi}(y|\mathbf{z}_{s})p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})}{p_{\phi}(y|\mathbf{z}_{t})}$ (47)
Since the model parameters implicitly learn the normalized distribution, we can drop the $p_{\phi}(y|\mathbf{z}_{t})$ term. Then, at each iteration, we update the parameters $\theta,\phi$ in the direction of the gradient of $\log p_{\theta,\phi}(\mathbf{z}_{s}|\mathbf{z}_{t},y)$ obtained as the sum of the gradients of the unconditional distribution $\log p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$ and classifier probability $p_{\phi}(y|\mathbf{z}_{s})$ with respect to the sampled sequence $\mathbf{z}_{s}$.
@@
Unlike classifier-free guidance, classifier-based guidance does not require training a generative model from a conditioned dataset. However, the gradient-based strategy for classifier-based guidance is not directly applicable to discrete state spaces due to the lack of a defined gradient. To mimic gradient-based updates to each sampling step, Gruver et al. [87] leveraged iterative guidance steps on continuous latent embeddings for each token before decoding back to a discrete sequence at each time step. However, projecting to and from the continuous and discrete spaces results in inconsistencies in the guided diffusion process, where optimized hidden embeddings do not always map to optimal tokens.
@@
However, computing $p_{\phi}(y|\mathbf{z}^{\prime}_{s})$ for all the possible transitions from state $\mathbf{z}_{t}$ is computationally intractable. Previous work has bypassed this limitation by approximation. For continuous and differentiable classifier functions $p(y|\mathbf{x}):\mathbb{R}^{L\times|\mathcal{V}|}\to\mathbb{R}$, we can approximate the denominator using the first-order Taylor expansion given by
$\log p_{\phi}(y|\mathbf{z}_{s},\mathbf{z}_{t})$ $\approx\log p_{\phi}(y|\mathbf{z}_{t})+(\mathbf{z}_{s}-\mathbf{z}_{t})^{\top}\nabla_{\mathbf{z}}\log p(y|\mathbf{z})|_{\mathbf{z}=\mathbf{z}_{t}}$ (50)
which approximates the likelihood of observing property $y$ at the slightly denoised state $\mathbf{z}_{s}=\mathbf{z}_{t-\frac{1}{T}}$ given the known log-probability of observing $y$ for state $\mathbf{z}_{t}$. This eliminates the need to explicitly sample $\mathbf{z}_{s}$ for all possible state transitions and compute $\log p_{\phi}(y|\mathbf{z}_{s},\mathbf{z}_{t})$.
Digress [113] has used this approximation strategy for guided discrete diffusion on categorical graph generation. Furthermore, Nisonoff et al. [27] uses the Taylor-approximated conditional distribution $\log p_{\phi}(y|\mathbf{z}_{s})$ to adjust the unconditional transition rates $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s}|y)$ given the unconditional rates $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s})$ for predictor-guidance of Continuous-Time Markov Chains (CTMCs) in the discrete state space.
@@
where $R_{t}(\mathbf{z}_{t},\mathbf{z}_{s}|y)$ is the predictor-adjusted rate of transitioning from state $\mathbf{z}_{t}$ to state $\mathbf{z}_{s}$
However, this strategy fails to scale to multiple objective guidance since it would require computing the joint probability over $K$ objectives $p_{\phi}(y_{1},y_{2},\dots,y_{K}|\mathbf{z}_{s},\mathbf{z}_{t})$ for some $K>1$. If all properties are mutually independent, we can factorize the distribution and compute the estimated probability of each objective and take their product $p_{\phi}(y_{1},y_{2},\dots,y_{K}|\mathbf{z}_{s},\mathbf{z}_{t})=\prod_{k=1}^{K}p_{\phi}(y_{k}|\mathbf{z}_{s},\mathbf{z}_{t})$. For the majority of multi-objective tasks including therapeutic peptide generation, independence across properties is not a reasonable assumption, and computing the joint distribution is required. Moreover, for objectives that guide toward contradictory optimal rates or transitions, training a model conditioned on these objectives could prevent the model from generating optimal sequences for either objective. Given these limitations, there remains a gap for efficient classifier-based conditioning for discrete diffusion that is robust to multi-objective tasks, which we address in this work.
@@ B.1 State-Dependent Masking Schedule
From Equation (34), we define the continuous-time forward masking probability $1-\alpha_{t}$ at time $t$ with $\alpha_{t}=\exp(-\bar{\sigma}(t))$ , where $\bar{\sigma}:[0,1]\to\mathbb{R}^{+}$ is the cumulative discrete-time masking schedule. Following Lou et al. [114], we apply a log-linear masking schedule $\bar{\sigma}(t)=-\log(1-t)$ for the forward diffusion process which has shown to result in the lowest variance in the NELBO loss [104]. Therefore, the continuous-time probability of remaining unmasked at time $t$ is equal to $\alpha_{t}=\exp\big{(}-(-\log(1-t))\big{)}=1-t$ and the weight that scales the negative log loss (NLL) is given by $\frac{1}{t}$ by our derivation in Appendix B.3.
For peptide-bond tokens, we alter the masking schedule such that peptide-bonds are masked at a slower rate at earlier time steps by defining a log-polynomial masking schedule $\bar{\sigma}(t)=-\log(1-t^{w})$, for some constant exponent $w>1$. Note that when $w=1$, the log-polynomial schedule reduces to the log-linear schedule. Therefore, the probability of remaining unmasked becomes $\alpha_{t}=\big{(}-(-\log(1-t^{w}))\big{)}=1-t^{w}$ and the weight that scales the negative log loss (NLL) is given by $\frac{w}{t}$ by our derivation in Appendix B.3.
@@
The NLL of the peptide-bond tokens is weighted more heavily than non-peptide bond tokens for $t$ close to 1. As $t\to 0$, the NLL weight approaches $\infty$ for all tokens. This biases the reverse diffusion process to unmask peptide bond tokens earlier since it was trained to minimize the loss associated with each unmasking step. As $t\to 0$, the large NLL weight ensures that the final unmasking steps during the reverse diffusion process result in an unmasked sequence that lies within the space of valid peptide SMILES.
Figure 15: Plots of state-dependent masking schedules. (A) The probability of remaining unmasked during the continuous-time forward diffusion process over time $t$ given different values of $w$ as the exponent of the masking schedule $\alpha_{t}=1-t^{w}$. We use $w=1$ for non-peptide bond tokens and $w=3$ for peptide bond tokens, resulting in slower masking of peptide-bond tokens. (B) The weight of the negative log-loss for different exponents $w$ in the log-polynomial masking schedule. The weight of the loss is higher for larger $w$ in earlier time steps, which results in a higher penalty for inaccurate predictions of peptide bond tokens compared to other tokens.
@@ B.2 Derivation of State-Dependent Reverse Posterior
For a single token, the state-dependent forward diffusion process is defined by the probability distribution $q(\mathbf{z}_{t}|\mathbf{x}_{0})$ which transforms the clean inputs to sequences with varying degrees of masking based on a probability distribution $\alpha_{t}(\mathbf{x}_{0})$. We define $\alpha_{t}(\mathbf{x}_{0}):\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}$ as a function that takes the clean token encoding $\mathbf{x}_{0}$ and outputs the probability of remaining unmasked at time $t$ depending on whether $\mathbf{x}_{0}$ encodes a peptide bond token.
@@
In this work, we classify each token into one of two states: peptide-bond tokens and non-peptide-bond tokens, which represent amino acid side chains and modifications. We define a function that generates a mask with values of 1 indicating tokens containing or contained within a peptide bond, and 0 otherwise (Algorithm 7). Let $\mathbf{b}\in\mathbb{R}^{|\mathcal{V}|}$ denote a vector with ones at indices of peptide-bond tokens. For derivation purposes, we let $\mathbf{b}^{\top}\mathbf{x}_{0}^{(\ell)}=1$ and $\mathbf{x}_{0}^{(\ell)}=\mathbf{b}$ when a token at position $\ell$ is a peptide bond token. Note that $\mathbf{b}$ is defined differently depending on the context of the token in the full sequence which is handled by the BondMask function. Then, we have $\alpha_{t}(\mathbf{x}_{0})=\big{(}\mathbf{1}-\mathbf{b}^{\top}\mathbf{x}_{0}\big{)}(1-t)+\mathbf{b}^{\top}\mathbf{x}_{0}(1-t^{w})$ or equivalently we can write
@@
This demonstrates that the probability of remaining in a masked state when $\mathbf{x}_{0}=\mathbf{b}$ is smaller than when $\mathbf{x}_{0}\neq\mathbf{b}$ since taking the exponent of a fraction results in a smaller value. So we have $\frac{s^{w}}{t^{w}}<\frac{s}{t}$ for $w>1$.
@@ B.3 Derivation of State-Dependent NELBO Loss
$\mathcal{L}_{\text{NELBO}}$ $=\sum_{n=1}^{T-1}\mathbb{E}_{q(\mathbf{z}_{t(n)}|\mathbf{x}_{0})}\bigg{[}\text{KL}\bigg{(}q(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)},\mathbf{x}_{0})||p_{\theta}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})\bigg{)}\bigg{]}$
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}\bigg{[}T\cdot\text{KL}\bigg{(}q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})\big{|}\big{|}p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})\bigg{)}\bigg{]}$ (68)
@@
where $o\left(\frac{1}{T^{2}}\right)$ denotes higher order terms that grow slower than $\frac{1}{T^{2}}$.
@@
$\mathcal{L}_{\text{NELBO}}=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}\bigg{[}T\cdot\text{KL}\bigg{(}q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})\big{|}\big{|}p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})\bigg{)}\bigg{]}$
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}T\cdot\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{w}{tT}-o\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$
@@
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{wT}{tT}-To\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$
@@
$=\mathbb{E}_{t\in\{\frac{1}{T},\frac{2}{T},\dots,1\}}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}\bigg{[}-\mathbf{1}[\mathbf{x}_{0}=\mathbf{b}]\left(\frac{w}{t}-To\left(\frac{1}{T^{2}}\right)\right)\log\langle\mathbf{x}_{0},\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\rangle$